OpenStudy (anonymous):

how does (1-sin(x)/cos(x))/(sin(x)cos(x))=(cos(x)-sin(x))/(cos(x)(sin(x)-cos(x))) ?

OpenStudy (anonymous):

\[\frac{1-\sin x \div \cos x }{\sin x-\cos x}=\frac{\cos x -\sin x}{\cos x(\sin x -\cos x)}\]

OpenStudy (noelgreco):

Multiply numerator and denominator by cos x

OpenStudy (anonymous):

*facepalm* thanks